How we implemented an LDAP directory for Laboratories

A Case Study at Hong Kong Institute of Vocational Education (Tsing Yi), Department of ICT

   Nick Urbanik <nicku@nicku.org>

Copyright Conditions: Open Publication License (see

  http://www.opencontent.org/openpub/  )

Sydney Linux Users Group (SLUG) Building 2, Level 4, Room 410, UTS Broadway

24 June 2005, 8.20 pm

July 10, 2005

Nick Urbanik <nicku@nicku.org> Copyright Conditions: Open Publication License (see http://www.opencontent.org/openpub/ ) Sydney Linux Users Group (SLUG) Building 2, Level 4, Room 410, UTS Broadway 24 June 2005, 8.20 pm

How we implemented an LDAP directory for Laboratories

What do you already know about ldap? How many know that a directory is tree shaped? How many have worked with a directory before? How many know about snmp object ids? How many know … What… is the air-speed velocity of an unladen swallow? The European swallow appears to do about 11 ms^{−1}

How we implemented an LDAP directory for Laboratories

What Do You Want? I could talk for ten hours (actually, I could go on for twenty after a few beers)

… but we have only an hour

What topics do you want us to cover here?

I think we need to understand the basics of ldap itself

operations some simple tools

… to make sense of other topics, especially programming The basics of the way inheritance works in ldap is important when understanding how to design a schema

How we implemented an LDAP directory for Laboratories

[bm=Argument for LDAP,toc=Argument for LDAP]Reasons for ldap and problems with alternatives

We don’t have time for this!

How we implemented an LDAP directory for Laboratories

Account Information The computer uses numbers to refer to users and groups Humans prefer to use names (like nicku) When you create files in your shared network drive, the client must access them using the same numbers The user ID numbers and group ID numbers must be the same on all computers Otherwise won’t be able to read own files!

How we implemented an LDAP directory for Laboratories